# Gaming on Coincident Peak Shaving: Equilibrium and Strategic Behavior

### v4

@@ VI-A Multi-agent CP game
In the Proposition, we present the condition for $S_{\mathrm{b},1}<S_{\mathrm{b},2}$; when $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$, the only difference is to replace ±$\pm r_{i}$ with $\mp r_{i}$. This Proposition demonstrates that we can analytically determine the equilibrium solution for one set of agents — either the CP-period agent set $N_{\mathrm{cp}}$ or non-CP-period agent set $N_{\mathrm{ncp}}$. For the remaining set, we can determine only the aggregated performance of the set. Thus, we call it a hybrid NE, where one subset of agents follows a strict NE, while another subset allows internal flexibility while maintaining the set-level balance. Essentially, more flexible agents, those with lower shifting penalty parameters and higher baseline demand differences, will shift more to achieve the overall shifting performance of their set. One important insight from this proposition is that system demand remains balanced under non-concave game conditions, aligning with the two-agent setting analysis. This result helps in evaluating system performance in terms of PoA and peak shaving effectiveness.
@@
We then analyze the stability of the equilibrium point in the multi-agent CP game $G^{\prime}$ based on the two-agent game results presented in Theorem 6. Extending the gradient vector to include all agents’ shifting decision variables $x_{i},i\in N$, we still denote the system dynamics as (9). Following the same proof structure, we can derive the multiple Lyapunov functions for periods 1 and 2, demonstrating that the global uniform asymptotically stable property also holds for the multi-agent CP game $G^{\prime}$. Under concave and quasiconcave conditions, the unique NE described in (22) and Proposition 12 serve as the stable equilibrium point in the system dynamics. In the non-concave condition, the system dynamics in periods 1 and 2 converge to the switching surface where the payoffs of both periods are equal, ensuring system demand remains balanced, as characterized by the hybrid NE in Proposition 13. The strategy set $\mathcal{X}_{\mathrm{s}}$ change to include all agents’ decisions as follows
@@
With this stability property, we then show that a finite learning rate can still be chosen such that a gradient-based algorithm using (12) as an updating rule finds the equilibrium point for the multi-agent CP game $G^{\prime}$. The only difference in the multi-agent setting is that the criteria shift from the convergence of each agent’s payoff across two periods to the convergence of the aggregated payoff of all CP-period agents and non-CP-period agents across the two periods. This also demonstrates that, under non-concave conditions, agents can internally adjust their demand-shifting strategies while ensuring that system demand remains balanced across the two periods.
@@
Determination of equilibrium point in multi-agent CP game. A finite learning rate vector $\tau_{h}\in\mathbb{R}^{N}$ of a finite difference approximation to a system dynamics in the multi-agent CP game setting can be chosen such that when $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$,
@@
$\sum_{i\in N_{\mathrm{ncp}}}-f_{i,2}(x_{i,\overline{h}})<\sum_{i\in N_{\mathrm{ncp}}}-f_{i,2}(x_{i,\underline{h}}),\sum_{i\in N_{\mathrm{ncp}}}-f_{i,1}(x_{i,\overline{h}})>\sum_{i\in N_{\mathrm{ncp}}}-f_{i,1}(x_{i,\underline{h}}).$ (26b)
When $S_{\mathrm{b},1}<S_{\mathrm{b},2}$, Greater than sign and less than sign reverses.
To conclude, the game framework is also workable for the multi-agent two-period setting as the (hybrid) NE still exists and is global uniform asymptotically stable within a strategy set as defined in (25). Also, a gradient-based algorithm with the updating rule that approximates the system dynamics can still compute the equilibrium point. We then analyze the impact of consumers’ strategic behavior in the multi-agent setting.
VI-B Consumers strategic behavior in Multi-agent CP game
We show the peak shaving effectiveness and PoA under the multi-agent setting in this section, where the centralized model could be denoted as $(x_{i,\mathrm{cen}}^{*},x_{-i,\mathrm{cen}}^{*})\in\arg\max_{x_{i},x_{-i}}\sum_{i\in N}-f_{i}(x_{i},x_{-i})$. Combined with Proposition 8, under the concave game condition, the centralized solution can be easily obtained by the first-order optimality conditions.
The peak shaving effectiveness, as defined in Theorem 9 with the extension from agent y to $-i$ denoting all agents but $i$, is still equal to 1 at the equilibrium of the game model. The reason is that the system demand will be balanced under all game conditions, the same as the centralized model. The PoA in the multi-agent setting is $P_{N}=\sum_{i\in N}f_{i}(x_{i}^{*},x_{-i}^{*})/\sum_{i\in N}f_{i}(x_{\mathrm{cen},i}^{*},x_{\mathrm{cen},-i}^{*})$. As the centralized model is convex from Proposition 8, and $x_{\mathrm{cen},i}^{*},x_{\mathrm{cen},-i}^{*}$ is the unique minimizer for the problem, we always have $P\geq 1$. Specifically, the concave game model remains equivalent to the corresponding centralized model and always has $P=1$.
With the agent number increase, the PoA under quasiconcave and non-concave games will be affected due to the game type changes, and we have the following remark.
@@ VI-B Consumers strategic behavior in Multi-agent CP game
It is intuitive that the non-concave game type is more likely to appear as the agent number increases because the probability that all agents are capable or non-capable reduces exponentially. As we analyzed at Theorem 11, the agent’s flexibility changes game types and thus causes different PoA. This means the PoA of a small system is more sensitive to the agents’ flexibility, and a large system can diminishing flexible agents influence on the system efficiency. This suggests a strategy for operators and utilities when implementing a game-based CP shaving program: selecting more flexible customers to form a large CP shaving system while grouping a few inflexible agents into a smaller CP shaving system. These two systems would operate separately, with individual CP periods determined by the customers within each system. This approach enhances overall system efficiency.
@@ VII Numerical Example
In this section, we use numerical simulations to show the CP game solution, and we show the numerical test aligns with our theoretical analysis. We set the CP charge price to $\pi=1$ and set three two-agent two-period cases and one multi-agent two-period case to show the CP game performance. In the multi-agent setting, we also test the influence of agents’ numbers on PoA.
VII-A Two-agent two-period CP game
@@
We set the baseline demand as $X_{1}=6,X_{2}=3,Y_{1}=3,Y_{2}=10$, then change the shifting penalty parameters to change the agent capability.
@@
(1) Set $\alpha_{\mathrm{x}}=0.2,\alpha_{\mathrm{y}}=0.1$, then all agents’ are capable according to Definition 3 and the game is quasiconcave. We show the solution path in Fig. 1, where the system alternates between periods 1 and 2. With each transition back to the previous period, the cost either decreases or increases in a way that reduces the cost difference for each agent between the two periods. Finally, the cost distance reduces to zero for each agent as they balance their demand over the two periods, and the system converges to the equilibrium point, corresponding to (7c). Compared with the solution from the centralized model, we observe that there are huge shifting changes and an increase in the system’s overall cost due to anarchy, reflected as the PoA is 1.125, indicating anarchy increases the system cost by 12.5% compared to the centralized method. The peak shaving ratio is the same as they all balance system demand.
Figure 1: Convergence of two-agent quasiconcave CP game.
@@
(2) Set $\alpha_{\mathrm{x}}=0.5,\alpha_{\mathrm{y}}=0.1$, then agent x is non-capable, and agent y is capable according to Definition 3, and the game is non-concave. We show the solution trajectory in Fig. 2, which is similar to the quasiconcave condition. However, each agent’s cost difference over the two periods does not reach zero upon convergence, as neither agent fully self-balances its demand. Instead, the more flexible agent (agent y) achieves greater cost reduction by lowering its demand during the CP period. The results align with (7d). Compared with the quasiconcave game, we observe that the demand shifting from the game model is closer to the centralized solution, and the system cost increase due to anarchy is also reduced with a PoA of 1.0941.
Figure 2: Convergence of two-agent non-concave CP game.
@@
(3) Set $\alpha_{\mathrm{x}}=0.5,\alpha_{\mathrm{y}}=0.6$, then both agents can’t change the CP period together, and the game is concave, where only period 2 is active during the entire solution trajectory. Thus, as we show in Fig. 3, the trajectory doesn’t jump and gradually reduces to the critical point, which corresponds to (7b). Obviously, the demand shifting converges to the same point with the centralized model, and the PoA is 1, indicating these two models are exactly equivalent.
Figure 3: Convergence of two-agent concave CP game.
VII-B Multiple agents and two-period CP game
@@
Fig. 4 shows the convergence and cost trajectory, and Fig. 5 shows the cost trajectory of non-CP period agent set and CP period agent set. Noted that the results align with our analysis in Proposition 13. Specifically, non-CP period agents shift demand either to the balance point or until they reach their critical limit. For example, agents 1 and 6 balance their demand across the two periods, while agent 3 reaches its shifting limit. After approximately 120 iterations, system demand balances over two periods (Figure 5), and CP-period agents begin internally adjusting their demand. Notably, agents 4 and 5, which have higher shifting penalty parameters than agent 2, gradually reduce their demand shifting, while agent 2 increases its shifting. From the cost trajectory (Figure 4), agent 4, with the highest shifting penalty, reduces its demand shifting more significantly than agent 5. Compared to the centralized model, the CP game equilibrium deviates significantly, with all agents shifting more demand due to the information barrier. This increases the cost and results in a PoA of 1.1317 while maintaining the same peak shaving performance.
TABLE I: Agents’ parameters and solution of game and centralized model
@@
PoA 1.1317
@@
Figure 4: Convergence of multi-agent concave CP game.
Figure 5: Convergence of CP period agent and non-CP period agent.
VII-C Agent number impacts on PoA
@@
In this section, we randomly generate agent samples for the systems with different agent numbers while satisfying the non-concave and quasiconcave game conditions as described in (20b) and (20c), i.e., $\sum_{i\in N}-r_{i}<b<\sum_{i\in N}r_{i}$. We set the agent’s $i,i\in N$ baseline demand as $X_{i,1},X_{i,2}\in(0,15)$ and penalty parameters as $\alpha_{i}\in(0,0.5)$. We loop the agent number from 2 to 50, and each agent number generates 1000 samples to calculate the PoA. We present the results in Fig. 6. As our theoretical analysis indicates, PoA is more fluctuates when the agent number is small; also, the game type is more varied with quasiconcave, concave, and non-concave, affected by agents’ flexibility, indicating the small system is more sensitive to the agent’s flexibility. When the agent number increases, all game conditions become non-concave, the mean PoA converges, and the variance decreases. This means the large system is more stable and can eliminate the agent’s flexibility influence on PoA.
@@
Figure 6: Agent number impacts on PoA. Each box contains 1000 samples, the circle denotes the outlier, the box upper and lower line denotes the quartiles, and the short blue line inside the box is the median value.

### v5

@@ VI-A Multi-agent CP shaving game
This proposition establishes that the equilibrium solution can be analytically determined for one group of agents, either the CP-period agent set $N_{\mathrm{cp}}$ or the non-CP-period agent set $N_{\mathrm{ncp}}$. However, for the remaining set, only the aggregate performance can be determined, rather than individual agent strategies. As a result, we refer to this as a hybrid NE, where one subset of agents follows a strict NE while the other retains internal flexibility, adjusting their individual strategies while maintaining the overall balance at the set level. In practice, more flexible agents, those with lower shifting penalties and greater baseline demand differences, will contribute more to the overall shifting performance of their set. A key insight from this result is that system demand remains balanced under non-concave game conditions, consistent with the findings from the two-agent setting. This conclusion provides a foundation for evaluating system efficiency losses and peak shaving effectiveness.
@@
We then analyze the stability of the equilibrium point in the multi-agent CP shaving game $G^{\prime}$ by building on the two-agent game results presented in Theorem 2. Extending the gradient vector to incorporate all agents’ shifting decision variables $x_{i},i\in N$, we still denote the system dynamics as (9). we continue to represent the system dynamics using (9). Following the same proof structure, we derive multiple Lyapunov functions for periods 1 and 2, demonstrating that the globally uniform asymptotically stability also holds for the multi-agent CP shaving game $G^{\prime}$. Under concave and quasiconcave conditions, the unique NE described in (22) and Proposition 3 serve as the stable equilibrium point in the system dynamics. In the non-concave case, the system dynamics in periods 1 and 2 converge to a switching surface where the payoffs of both periods are equal, ensuring system demand remains balanced, as characterized by the hybrid NE in Proposition 4. The strategy set $\mathcal{X}_{\mathrm{s}}$ is then extended to incorporate all agents’ decisions as follows:
@@
With this stability property, we further show that a finite learning rate can be selected such that a gradient-based algorithm using (12) as an updating rule converges to the equilibrium point in the multi-agent CP shaving game $G^{\prime}$. The key distinction in the multi-agent setting is that convergence shifts from each agent’s individual payoff across the two periods to the aggregated payoff of all CP-period agents and non-CP-period agents. This also confirms that under non-concave conditions, agents can internally adjust their demand-shifting strategies while maintaining overall system demand balance across the two periods.
@@
Determination of equilibrium point in the multi-agent CP shaving game. A finite learning rate vector $\tau_{\mathrm{cp},h}\in\mathbb{R}^{N_{\mathrm{cp}}}$ for CP-period agents in the finite difference approximation to the system dynamics in the multi-agent CP shaving game can be selected such that when $\sum_{i\in N_{\mathrm{cp}}}X_{i,1}+x_{i}\geq\sum_{i\in N_{\mathrm{cp}}}X_{i,2}-x_{i}$,
@@
When $\sum_{i\in N_{\mathrm{cp}}}X_{i,1}+x_{i}<\sum_{i\in N_{\mathrm{cp}}}X_{i,2}-x_{i}$,
$\sum_{i\in N_{\mathrm{cp}}}-f_{i,1}(x_{i,\overline{h}})>\sum_{i\in N_{\mathrm{cp}}}-f_{i,1}(x_{i,\underline{h}}),\sum_{i\in N_{\mathrm{cp}}}-f_{i,2}(x_{i,\overline{h}})<\sum_{i\in N_{\mathrm{cp}}}-f_{i,2}(x_{i,\underline{h}}).$ (26b)
The same also holds for the learning rate vector $\tau_{\mathrm{ncp},h}\in\mathbb{R}^{N_{\mathrm{ncp}}}$ for non-CP-period agents.
To conclude, the game framework remains effective in the multi-agent, two-period setting, as the (hybrid) NE continues to exist and exhibits globally uniform asymptotically stable within the strategy set defined in (25). Additionally, a gradient-based algorithm, using an updating rule that approximates the system dynamics of the multi-agent CP shaving game, can reliably compute the equilibrium point. Building on this foundation, we next analyze the impact of customers’ strategic behavior in the multi-agent setting.
VI-B Customers strategic behavior in multi-agent CP shaving game
In this section, we analyze the peak shaving effectiveness and efficiency loss in the multi-agent setting. The centralized model is formulated as $(x_{i,\mathrm{cen}}^{*},x_{-i,\mathrm{cen}}^{*})\in\arg\max_{x_{i},x_{-i}}\sum_{i\in N}f_{i}(x_{i},x_{-i})$, representing the optimal solution that maximizes the total agent payoffs. Combined with Proposition 2, under the concave game condition, the solution to the concave centralized model can be directly obtained using the first-order optimality conditions.
The peak shaving effectiveness, as defined in Theorem 4, extends from the two-agent setting to the multi-agent setting by replacing agent y to $-i$, denoting all agents except $i$. It remains equal to 1 at the equilibrium of the game model because system demand is always balanced under all game conditions, just as in the centralized model. The efficiency loss in the multi-agent setting is given by $P_{N}=\sum_{i\in N}f_{i}(x_{i}^{*},x_{-i}^{*})/\sum_{i\in N}f_{i}(x_{\mathrm{cen},i}^{*},x_{\mathrm{cen},-i}^{*})$. Since the centralized model is concave, derived from Proposition 2, and $x_{\mathrm{cen},i}^{*},x_{\mathrm{cen},-i}^{*}$ is its unique minimizer, we always have $P\geq 1$. Specifically, under the concave game condition, the game model is equivalent to the centralized model, ensuring that $P=1$.
As the number of agents increases, the efficiency loss under quasiconcave and non-concave games is influenced by changes in the game structure. This leads to the following remark.
@@ VI-B Customers strategic behavior in multi-agent CP shaving game
It is intuitive that the non-concave game type becomes more prevalent as the number of agents increases because the probability that all agents are either capable or non-capable decreases exponentially. As analyzed in Theorem 6, agents’ flexibility influences the game structure and, consequently, the efficiency loss. This implies that smaller systems are more sensitive to variations in agent flexibility, whereas larger systems can mitigate the impact of highly flexible agents on overall efficiency.
This insight provides guidance for operators and utilities implementing game-based CP shaving programs. Specifically, selecting more flexible customers to form a larger CP shaving system while grouping a few inflexible agents into a smaller CP shaving system can enhance efficiency. These two systems would operate independently, each determining its own CP periods based on the participating customers. This targeted structuring can improve overall system performance and mitigate efficiency loss.
@@ VII Numerical Example
In this section, we use numerical simulations to show the CP shaving game solution, and we show that the numerical test aligns with our theoretical analysis. We set the CP charge price to $\pi=1$ and set three two-agent two-period cases and one multi-agent two-period case to show the CP shaving game performance. In the multi-agent setting, we also test the influence of the number of agents on efficiency loss.
VII-A Two-agent two-period CP shaving game
@@
We set the baseline demand as $X_{1}=3,X_{2}=10,Y_{1}=6,Y_{2}=3$, then change the shifting penalty parameters to change the agent capability.
@@
(1) Set $\alpha_{\mathrm{x}}=0.1,\alpha_{\mathrm{y}}=0.2$, then all agents’ are capable according to Definition 3 and the game is quasiconcave. We show the solution path in Fig. 1, where the system alternates between periods 1 and 2. With each transition back to the previous period, the cost either decreases or increases in a way that reduces the cost difference for each agent between the two periods. Finally, the cost distance reduces to zero for each agent as they balance their demand over the two periods, and the system converges to the equilibrium point, corresponding to (7b). Compared with the solution from the centralized model, we observe that there are huge shifting changes and an increase in the system’s overall cost due to anarchy, reflected as the efficiency loss is 1.125, indicating anarchy increases the system cost by 12.5% compared to the centralized method. The peak shaving ratio is the same because they all balance the system demand.
Figure 1: Convergence performance of two-agent quasiconcave CP shaving game.
@@
(2) Set $\alpha_{\mathrm{x}}=0.1,\alpha_{\mathrm{y}}=0.5$, then agent y is non-capable, and agent x is capable according to Definition 3, and the game is non-concave. We show the solution trajectory in Fig. 2, which is similar to the quasiconcave condition. However, each agent’s cost difference over the two periods does not reach zero upon convergence, as neither agent fully self-balances its demand. Instead, the more flexible agent (agent x) achieves greater cost reduction by lowering its demand during the CP period. The results align with (7c). Compared to the quasiconcave game, we observe that the demand shifting resulting from the game model more closely approximates the centralized solution, and the increase in system cost due to anarchy is reduced, with an efficiency loss of 1.0941.
Figure 2: Convergence of two-agent non-concave CP shaving game.
@@
(3) Set $\alpha_{\mathrm{x}}=0.6,\alpha_{\mathrm{y}}=0.5$, then both agents can’t change the CP period together, and the game is concave, where only period 2 is active during the entire solution trajectory. Thus, as we show in Fig. 3, the trajectory doesn’t jump and gradually reduces to the critical point, which corresponds to (7a). Obviously, the demand shifting converges to the same point as the centralized model, and the efficiency loss is 1, indicating these two models are exactly equivalent.
Figure 3: Convergence performance of two-agent concave CP shaving game.
VII-B Multiple agents and two-period CP shaving game
@@
Fig. 4(a) shows the convergence and cost trajectory, and Fig. 4(b) shows the cost trajectory of the non-CP period agent set and the CP period agent set. Noted that the results align with our analysis in Proposition 4. Specifically, non-CP period agents shift demand either to the balance point or until they reach their critical limit. For example, agents 1 and 6 balance their demand across the two periods, while agent 3 reaches its shifting limit. After approximately 120 iterations, system demand balances over two periods (Fig. 4(b)), and CP-period agents begin internally adjusting their demand. Note that agents 4 and 5, which have higher shifting penalty parameters than agent 2, gradually reduce their demand shifting, while agent 2 increases its shifting. From the cost trajectory (Fig. 4(a)), agent 4, with the highest shifting penalty, reduces its demand shifting more significantly than agent 5. Compared to the centralized model, the CP shaving game equilibrium deviates significantly, with all agents shifting more demand due to the information barrier. This increases the cost and results in an efficiency loss of 1.1317 while maintaining the same peak shaving performance.
TABLE I: Agents’ parameters and solution of the game and centralized model
@@
Efficiency loss 1.1317
(a) Individual agent
(b) CP period and non-CP period agent
@@
Figure 4: Convergence performance of multi-agent non-concave CP shaving game.
VII-C Agent number impacts on efficiency loss
@@
In this section, we randomly generate agent samples for systems with varying agent numbers while ensuring that the generated samples satisfy the non-concave and quasiconcave game conditions described in (20b) and (20c). We allow the baseline system demand consumption to vary freely, thus getting the range $\sum_{i\in N}-r_{i}<b<\sum_{i\in N}r_{i}$. We set the agent’s $i,i\in N$ baseline demand as $X_{i,1},X_{i,2}\in(0,15)$ and penalty parameters as $\alpha_{i}\in(0,0.5)$. We loop the agent number from 2 to 50, and each agent number generates 1000 samples to calculate the efficiency loss. We present the results in Fig. 5. As our theoretical analysis indicates, efficiency loss is more variable when the agent number is small; also, the game type is more varied with quasiconcave, concave, and non-concave, affected by agents’ flexibility, indicating that the small system is more sensitive to the agents’ flexibility. When the agent number increases, all game conditions become non-concave, the mean efficiency loss converges, and the variance decreases. This means the large system is more stable and can eliminate the agent’s flexibility influence on efficiency loss.
@@
Figure 5: Agent number impacts on efficiency loss. Each box contains 1000 samples, the circle denotes the outlier, the box upper and lower line denotes the quartiles, and the short blue line inside the box is the median value.
VII-D Real-world applications
In this section, we apply our CP shaving game model to a real-world case in Texas. We utilize data from ERCOT’s 4CP program, under which customers pay a coincident peak charge based on their average demand during one system peak hour in each of the four summer months (June through September). The 4CP charge rate is generally determined by the transmission cost of service (TCOS), which was approximately \$66.76 per kilowatt-year in 2024, implying that a 1 kW peak demand results in \$66.76 in annual electricity bill [Kavulla,, 2025]. ERCOT publicly provides each participant’s CP demand record along with the exact CP hours [ERCOT, 2025a, ]. In this study, we focus on the year 2024, during which the CP hours occurred on June 30 (17:45), July 1 (17:00), August 20 (17:00), and September 20 (16:00), and the total participants is 136 (excluding 6 entities with zero CP demand). The final CP demand is the average of these four hourly CP demands.
Due to the proprietary nature of detailed consumption data, and given that our primary focus is evaluating CP demand reduction under a game-theoretic framework, we aggregate all hours from June 1 to September 30 outside the four CP intervals into a single non-CP period. We then use the average demand across this aggregated set to represent each participant’s off-peak consumption. Using ERCOT’s system-wide hourly demand data [ERCOT, 2025b, ], we estimate the average non-CP period demand for each participant. This approach is based on the assumption that large commercial and industrial customers—the primary participants in the 4CP program—have electricity consumption patterns that are largely driven by their operational capacity and schedules, and typically exhibit relatively consistent load shapes throughout the day [Jardini et al.,, 2000; Cropp et al.,, 2021]. Although some temporal fluctuations exist, averaging over a multi-period window helps smooth out short-term volatility. Based on this, we estimate each participant’s average non-CP period demand by applying its demand ratio in the CP period to the system-wide average non-CP period demand, with an added 50% random variation to capture heterogeneity across customers.
We then determine each company’s shifting cost parameter based on its demand value. Specifically, we calculate a minimum shifting cost parameter $\alpha_{i}^{\prime}$ using the period 2 demand, which, under our assumption, is the baseline CP period. This ensures that the customer’s demand does not become negative during the game, and is given by the expression $\alpha_{i}^{\prime}=\pi/2X_{i,2}$. To capture demand flexibility and introduce heterogeneity into the simulation, we add a 20% random fluctuation to the minimum cost parameter.
Fig. 6(a) presents the CP charges based on actual CP demand data and the charges after applying the CP shaving game model. Markers are used to distinguish CP-period and non-CP-period customers. We can see that customers with higher baseline demand tend to bear a larger share of CP charges in the real system. Because their consumption has a greater impact on the system peak, and because they generally face lower shifting costs, these customers tend to benefit more under the game framework. System-wide, the game leads to a significant reduction in CP charges compared to the real CP-influenced demand data—approximately \$651 million. This effect is further illustrated in the convergence behavior shown in Fig. 6(b), where the total cost (i.e., CP charge plus shifting cost) declines for CP-period customers and slightly increases for non-CP-period customers. The shape of the convergence curve also suggests that the game exhibits quasiconcavity, which is a result of the way we specify the shifting cost parameters. The observed reduction in CP charges arises because we assume all customers in the game model strategically participate to minimize their peak-related cost. In practice, not all customers are interested in investing in CP-shaving efforts. Therefore, our results suggest that increasing customer participation has the potential to further drive down the CP demand and improve system efficiency.
We also compare the game results with those from a centralized peak shaving model, and observe that both approaches always achieve the same peak shaving effectiveness. To evaluate efficiency loss, we vary the random variation in the demand ratio across five levels, 25%, 50%, 75%, 100%, and 125%, and generate 50 random samples for each level to assess robustness. We show the efficiency loss in Fig. 6(c), which increases with the magnitude of variation in non-CP demand relative to CP demand. This is because greater variation requires more interactions to reach convergence. Still, the efficiency loss is mainly within 1.05 and demonstrates robust performance. This aligns with our theoretical analysis and can be attributed to the stabilizing effect of a large number of participating agents.
(a) CP demand performance
(b) Convergence curve of the CP game
(c) Uncertainty impact on demand pattern and shifting cost parameter
Figure 6: Results with Texas 4CP program.